OpenStudy (anonymous):

What is the equation to this Alternating Sequence? a: 1 4 7 b:0 8 16

OpenStudy (anonymous):

I don't think they are alternating series; alternating series switch signs

OpenStudy (anonymous):

i think i may have put that wrong, i need to get from a to b somehow. i think that i have to do 2 2/3a - something, thats what i've got so far.

OpenStudy (tad1):

For an increase of 8 for b, there is an increase of 3 for a. So if we plotted a line, it would have the slope of 8/3. When a = 1, b= 0 so the line has the point (1, 0 ) on it. so we can write a linear equation using the point-slope form of the line:\[y-y1 = m(x-x1)\]